OpenStudy (anonymous):

Find the Length of the curve y = e^x 0<=x<=5

OpenStudy (anonymous):

\[y=e^x 0 \le x \le5\]

OpenStudy (anonymous):

there is suppose to be a space after e^x

OpenStudy (anonymous):

o to 5 is the limit

OpenStudy (anonymous):

they have suggest u sub twice for me. but that has just confused me

OpenStudy (anonymous):

\[\int\limits_{0}^{1} \sqrt{1+(e^ (2x))}\]

OpenStudy (anonymous):

that is e^2x i dont know how to make the x up there with the 2 on the equation thing

OpenStudy (kainui):

``` e^{2x} ``` That'll do it =)

OpenStudy (anonymous):

this is where they want me to do the first u sub

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

i guess the e has confused me

OpenStudy (kainui):

Sure, this is kind of a weird integral. What did they suggest to you that was confusing? It might be better to just explain what they're telling you to do, since they probably picked the best thing already.

OpenStudy (anonymous):

ok, i will type the equation for you. give me a min

OpenStudy (anonymous):

\[\int\limits_{1}^{e} \sqrt{\frac{ 1+u^2 }{ u^2 }}udu\]

OpenStudy (anonymous):

where u= e^x x= lnu then dx=du/u

OpenStudy (anonymous):

but Idk how they got that

OpenStudy (anonymous):

its like, they show me WHAT they got, but not HOW

OpenStudy (anonymous):

then, they go right into the 2nd u sub here

OpenStudy (anonymous):

v= \[\sqrt{1+u^2}\] v^2=\[1+u^2\] and vdv= udu

OpenStudy (anonymous):

then \[\int\limits_{\sqrt{2}}^{\sqrt{1+e^2}} \frac{ v }{ v^2-1 }vdv\]

OpenStudy (anonymous):

this is where I am COMPLETELY lost now! I cant figure out what they did

OpenStudy (anonymous):

why would the limits become what they did?

OpenStudy (kainui):

Ok, also I think you might be wrong when you wrote these down, I think they should have been \[\int\limits_{1}^{e} \sqrt{\frac{ 1+u^2 }{ u^2 }}du\] and then \[\int\limits_{\sqrt{2}}^{\sqrt{1+e^2}} \frac{ v }{ v^2-1 }dv\] instead, without the extra u and v next to du and dv.

OpenStudy (anonymous):

i that that was weird too

OpenStudy (kainui):

Wait, I'm also confused, I thought you said the limits were 0 to 5 and here you've written 0 to 1? \[\int\limits_{0}^{1} \sqrt{1+e^ {2x}}\]

OpenStudy (anonymous):

cuz they made du, du/u

OpenStudy (anonymous):

yes... im sorry... I meant 1! I am having a hard time with equation thing and I get confused

OpenStudy (anonymous):

I have a 2nd problem where it is five.

OpenStudy (anonymous):

this one is 1

OpenStudy (kainui):

That's fine, here I'll show you how I will do the first substitution step by step. =)

OpenStudy (anonymous):

thank you

OpenStudy (kainui):

Before I get started, this is the correct integral, right? \[\Large \int\limits_{x=0}^{x=1}\sqrt{1+e^{2x}}dx\]

OpenStudy (anonymous):

yes

OpenStudy (kainui):

First make the substitution\[\Large u = e^x \] Now take its derivative \[\Large \frac{du}{dx}=e^x\] Multiply both sides by dx \[\Large du = e^x dx\] Now we look and see we need to plug in e^x and dx from our original integral. So we need to solve for dx: \[\Large \frac{du}{e^x}=dx\] and we use the substitution to make this correct: \[\Large dx=\frac{du}{u}\] So we can plug these in: \[\Large \int\limits_{x=0}^{x=1} \sqrt{1+(u)^2} ( \frac{du}{u})\] but notice our bounds are still in terms of x! That's ok, we now just plug in our bounds to the substitution we made as well! \[\Large u=e^0=1 \\ \Large u=e^1=e\] just replace the bounds: \[\Large \int\limits\limits_{u=1}^{u=e} \frac{\sqrt{1+u^2}}{u}du \]We can of course square and square root the bottom to make it look similar to what they have: \[\Large \int\limits_{u=1}^{u=e}\sqrt{\frac{1+u^2}{u^2}}du\] and that's it for this part.

OpenStudy (anonymous):

ok. let me look at this for a min plz

OpenStudy (anonymous):

before we move on, could u tell me how the u becomes squared in the denominator?

OpenStudy (anonymous):

oh wait... i see

OpenStudy (anonymous):

because you took the root of it?

OpenStudy (anonymous):

now...how do the limits become \[\int\limits_{\sqrt{2}}^{\sqrt{1+e^2}}\]

OpenStudy (kainui):

Try to do the same thing I just did on your own with the new substitution and I'll help you out. \[\Large v=\sqrt{1+u^2}\]

OpenStudy (anonymous):

ok. because they are removing the square root and changing 1-e to 2-the function

OpenStudy (anonymous):

ok, I have the final answer, but I was wondering if i could see how someone did their arithmetic.

OpenStudy (anonymous):

i believe partial fractions was used

OpenStudy (kainui):

Yeah I think so too. Want to show me your work on attempting the partial fractions? It doesn't do either of us any good if I just do it, it's better for you to get some practice and I can help see where you need help. I know like two or three methods to do partial fractions so I might be able to give you a tip to make it easier for you.

OpenStudy (anonymous):

sure. reminder... I am terrible at equation thing and will need some time to put all of this in

OpenStudy (kainui):

Haha believe me when I say I wasn't born knowing how to do partial fractions either. I practiced messing them up tons of times XD

OpenStudy (anonymous):

\[\int\limits_{\sqrt{2}}^{\sqrt{1+e^2}} \frac{ v }{ v^2-1}dv\] =\[\int\limits_{\sqrt{2}}^{\sqrt{1+e^2}} 1+(\frac{ 1/2 }{ (v-1)(v+1) }dv\] =\[[v+\frac{ 1 }{ 2 }\ln(\frac{ v-1 }{ v+1 })]\]=\[(\sqrt{1+e^2}+\frac{ 1 }{ 2 }\ln (\frac{ \sqrt{1+e^2}-1 }{ \sqrt{1+e^2}+1 }))-(\sqrt{2}+\frac{ 1 }{ 2 }\ln(\sqrt{2}))\]

OpenStudy (anonymous):

final answer is \[\sqrt{1+e^2}-\sqrt{2}+\ln (\sqrt{1+e^2}-1)-1-\ln(\sqrt{2}-1)\]

OpenStudy (anonymous):

i still feel confused lol but that ok. you helped me a lot

OpenStudy (anonymous):

L= of course

OpenStudy (kainui):

Hmm, sorry I'm having trouble concentrating, I stayed up all night working on stuff so I guess I'm finally braindead haha. I'll take a look tomorrow if you still want me to.

OpenStudy (anonymous):

sure :) i truly appreciate the help i get. And..I understand. Somedays, im like "I GOT THIS" and the other days Its like I forget ALL THE RULES for EVERYTHING and don't know how to do it anymore

OpenStudy (anonymous):

especially with e and lne and all that dumb stuff like square roots! My brain gets confused just looking at it

OpenStudy (kainui):

Haha yeah I know what you mean. It's a good feeling when you finally understand, actually I think one of the things that helped me as to teach other people. Somehow explaining things to other people is almost like I'm explaining it to myself and I end up learning at the same time somehow haha.